A car preceding down a 1 in 10 incline at 6.5kph when it's brake are abruptly applied. How far does it skid along the road before stopping If the wheels lock and the coefficient of friction between tyres and the road is 0.7?

1
Expert's answer
2022-02-20T15:45:23-0500

"s=\frac{v^2}{2g(\sin\alpha-\mu\cos\alpha)}=32.7~m."